OpenStudy (anonymous):

Find the value of 2 times the square root of 3, all over the square root of 10. in simplest form. 2 times the square root of 30. 2 times the square root of 3, all over 5. 2 times the square root of 30, all over 10. the square root of 30, all over 5.

OpenStudy (anonymous):

2 sqrt3/sqrt10 = 2 sqrt (3/10) = 2 sqrt( 3*10/10*10) = 2 sqrt30/10 = (sqrt30)/5

OpenStudy (anonymous):

Can you explain this please?

OpenStudy (anonymous):

\[2\sqrt3/\sqrt10= 2\sqrt(3/10)\] now inside the sqrt sign multiply the denominator and numerator by 10 and then denominator is 10 square so take it outside

OpenStudy (anonymous):

OKay thak youuu